OpenStudy (anonymous):

Determine whether the vectors u and v are parallel, orthogonal, or neither. u = <10, 6>, v = <9, 5> Orthogonal Parallel Neither

OpenStudy (anonymous):

@Zekarias

OpenStudy (anonymous):

Can u find the dot product of the two vectors?

OpenStudy (anonymous):

u*v = 0 -> orthogonal u*v = |u| * |v| ->parallel

OpenStudy (anonymous):

6 and 9?

OpenStudy (anonymous):

What are 6 and 9?

OpenStudy (anonymous):

im confused. im not sure how to solve this, or find dot product.

OpenStudy (anonymous):

given u = (u1,u2) and v = (v1,v2) the dot product is given by : u*v = u1v1 + u2v2

OpenStudy (anonymous):

\[u.v =|u||v|\cos(\theta)\]Then find \theta\

OpenStudy (anonymous):

in our case : u*v = 10*9 + 6*5 = 90 + 30 = 120 now its not 0 then they are not orthogonal we have to check now if 120 = |u||v| |u| = sqrt(10^2 + 6^2) = sqrt(136) |v| = sqrt(9^2 + 5^2) = sqrt(106) 120 = 120.066 ? no .. so they are not parallel..

OpenStudy (anonymous):

so its neither!

OpenStudy (anonymous):

right

OpenStudy (anonymous):

thank you mind helping me with one more problem?

OpenStudy (anonymous):

ask

OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <6, -1>, v = <7, -4> 20.3° 10.2° 0.2° 30.3°

OpenStudy (anonymous):

so u*v = |u|*|v|*cos(a) we can calculate u*v using the coordinates : 6*7 + (-1)*(-4) = 46 now |u| = sqrt(6^2 + (-1)^2) = sqrt(37) |v| = sqrt(7^2 + (-4)^2) = sqrt(49 + 16) = sqrt(65) so we have : 46 = sqrt(37) * sqrt(65) * cos(a) cos(a) = 46 / (sqrt(37) * sqrt(65)) a = arccos( 46 / (sqrt(37) * sqrt(65))) a=20.28 is aprox a=20.3

OpenStudy (anonymous):

thank you. you have been very helpful!

OpenStudy (anonymous):

yw.. ill sleep now lol have a good day

OpenStudy (anonymous):

Express the complex number in trigonometric form. -6i 6(cos 0° + i sin 0°) 6(cos 270° + i sin 270°) 6(cos 180° + i sin 180°) 6(cos 90° + i sin 90°)